OpenStudy (anonymous):

−2x + 3y = 8, 2x + 2y = 7 solve the linear equation

OpenStudy (anonymous):

add them together -2x + 2x +3y + 2y = 8+7 5y = 15 y = 3 2x + 2(3) = 7 2x =1 x = 0.5

OpenStudy (anonymous):

thank you